# Almost real closed fields with real analytic structure

### v1

@@ Abstract.
analytic structure, subanalytic sets, quantifier elimination, Weierstrass systems, Henselian valued fields, almost real closed fields
@@ 1. Introduction
Since the work of Lojasiewic, Gabrielov, and Hironaka from the sixties, it has been known that subanalytic sets in ${\mathbb{R}}^{n}$ exhibit tame behaviour similar to the semi-algebraic setting [23, 16, 19, 18]. In more modern terms, the structure ${\mathbb{R}}$ expanded with function symbols for restricted analytic functions is o-minimal. The original works by Lojasiewic, Gabrielov, and Hironaka are rather long and difficult, and work by Denef–van den Dries [10] greatly simplifies the treatment of subanalytic sets through the use of Weierstrass division. Since then, this Weierstrass division has been the central tool for understanding and developing analytic structures on fields, see e.g.  [12, 25, 11, 8, 7, 6, 9, 22, 1]. A comprehensive treatment of fields with analytic structure is the foundational work by Cluckers–Lipshitz [5], which includes both the case of real closed fields and henselian valued fields with analytic structure.
In the current article we continue this line of research, by focussing on fields equipped with real analytic structure, as defined in [5]. Up to now, these structures have only been considered when the field is real closed, leading to o-minimal structures. However, the notion of a field with real analytic structure makes sense even without this assumption, and the aim of this article is to develop the theory of arbitrary fields equipped with real analytic structure. As it turns out, any such field comes with a canonical henselian valuation for which the resulting residue field is ${\mathbb{R}}$. In particular, such fields are almost real closed.
@@
Our first main result is that almost real closed fields equipped with ${\mathcal{B}}$-analytic structure exhibit tame behaviour. Of course, in this generality one cannot hope for the theory to be o-minimal or even weakly o-minimal, simply because the field is not necessarily real closed. Instead we show that such fields are $\omega$-h-minimal, a tameness notion for valued fields developed by Cluckers–Halupczok–Rideau [3]. As a consequence of $\omega$-h-minimality, one automatically obtains that definable functions satisfy the Jacobian property, cell decomposition results, and dimension theory. Furthermore, one also obtains bounds on rational points on transcendental definable curves in a suitable sense [2], see Example 5.1.1. This counting of rational points was in fact the original motivation for this project.
@@
The proof of the above theorem consists of two main components: a relative quantifier elimination result, and an analysis of terms in one variable. For the relative quantifier elimination, we follow classical methods by Denef–van den Dries [10] and Flenner [14]. This is carried out in Section 4. The term analysis requires more work, and is partially based on a similar analysis by Cluckers–Lipshitz–Robinson [7] for real closed fields. In more detail, our approach is to first analyse terms over the real closure, and then pull back the resulting data to the ground field. The trickiest part in this term analysis is a precise control of parameters, for which we develop extension of parameters for real Weierstrass systems. This term analysis is the subject of Section 3.
The notion of $\omega$-h-minimality is really a tameness notion relative to the leading term structure $\mathrm{RV}$. As such, the above theorem does not readily give any information about the induced structure on the residue field and the value group. Therefore, our second main result gives a precise description of the induced structure on the residue field and the value group when $K$ is equipped with its canonical valuation. We state the following as an exemplar result, but see Theorem 5.2.1 for more general Weierstrass systems, as defined in Definition 2.1.1. We work in a certain three-sorted language ${\mathcal{L}}_{{\rm ac},{\mathcal{B}}}$ which includes an angular component map ${\rm ac}$, see Section 4.2 for a precise description. Denote by ${\mathbb{R}}_{\mathrm{an}}$ the structure ${\mathbb{R}}$ in the ring language expanded by symbols for restricted analytic functions.
@@
Let ${\mathcal{B}}={\mathcal{A}}((\Omega))$ be the full Weierstrass system over some ordered abelian group $\Omega$, and let $K$ be an almost real closed field with ${\mathcal{B}}$-analytic structure. We consider $K$ as a valued field with its canonical valuation. Then
@@
From this relative quantifier elimination, we also obtain an Ax–Kochen–Ersov type result. If $K$ and $K^{\prime}$ are almost real closed fields considered with the canonical valuation, then the classical Ax–Kochen–Ersov theorem tells us that $K$ and $K^{\prime}$ are elementarily equivalent as valued fields if and only if their value groups are elementarily equivalent. Indeed, the residue fields are both equal to ${\mathbb{R}}$. The next result can be seen as an extension of this, where we equip $K$ and $K^{\prime}$ with ${\mathcal{B}}$-analytic structure. Denote by ${\mathcal{L}}_{\mathrm{oag}}$ the language of ordered abelian groups $\{0,+,<\}$.
@@
Let ${\mathcal{B}}$ be a real Weierstrass system which is strong and rich, and let $K$ and $K^{\prime}$ be almost real closed fields with ${\mathcal{B}}$-analytic structure. We consider $K$ and $K^{\prime}$ as valued fields with their canonical valuation, and denote their value groups by $G$ and $G^{\prime}$. Then
@@ 2.1. Weierstrass systems
We recall the notion of a real Weierstrass system from [5, Def. 3.1.1]. Our definition is in fact slightly different from [5], and the differences will be explained in more detail below.
@@
Consider the system of algebras with $B_{n,\alpha}=A_{n,\alpha}((\Omega))$ for each $\alpha,n$. This is a (rich and strong) real Weierstrass system [5, Rem. 3.13(iv)]. It is called the full Weierstrass system over $\Omega$ and denoted by ${\mathcal{A}}((\Omega))$.
Our definition differs in two ways from [5], so let us indicate the changes. First, axiom 1(e) is new and simply tells us that we can permute the variables of the elements of a Weierstrass system. This seems to be necessary to ensure that composition is well-defined and is also present in the analogous definition of separeted Weierstrass system [6, Def. A.1.1]. Second, in the Weierstrass division axiom the value $\delta\in{\mathbb{R}}_{>0}$ does not depend on $m$, the number of variables of $g$. This will be important for us when expanding Weierstrass systems in Section 3.2. The richness condition is new, and will be needed for us when embedding a field with real analytic structure into a Hahn series field.
The notion of a strong Weierstrass system is also present in [5, Def. 3.1.1] and will be needed in Section 3.4 when analysing terms on the real closure. It roughly allows one to write $f(x,1/x)=g(x)+(1/x)h(1/x)$.
It is important to note that all natural examples of real Weierstrass systems from [5] satisfy the axioms above, in particular they are also real Weierestrass systems according to our definition. Indeed, in the Weierstrass division axiom the value radius $\delta$ depends only on $f$, and not on $g$, as follows from [17, Thm.   II.1, p.  80]. Additionally, all of these examples satisfy the richness condition ((4)). Since our notion of a real Weierstrass system is more restrictive, all results of [5] about them still hold in our context.
@@
This is part of [5, Rem. 3.1.4(ii)] and follows by Weierstrass dividing $1$ by $U(x)$, noting that the latter is regular of degree zero (in any variable). ∎
The following lemma is a real version of the implicit function theorem from [11, Rem. 1.3(4)] and is proved in a similar manner.
@@
We follow the proof of [11, Rem. 1.3(4)]. By Weierstrass division, we may write
@@
The proof idea of the lemma below is classical, see e.g.  [25, Lem. 2.4]. In this more abstract context, some additional input from Lemma 2.1.6 is required, as well as some care about radii of convergence.
@@
Since $\sigma\colon B_{0}\to K$ is order-preserving, (1d) allows us to reduce to the case that $\lVert f\rVert=1$. Using axiom (1c) and compactness of $[-1,1]_{{\mathbb{R}}}$, it suffices to prove this lemma for $a\in[-\delta,\delta]_{K}$ for some $\delta\leq\alpha$ with $\delta\in{\mathbb{R}}_{>0}$. Now, there exist some $b,c\in{\mathbb{R}}$ such that $b<f_{0}(0)<c$. By Lemma 2.1.7, applied to $f(x)-b$ and $f(x)-c$, it follows that also $b<\sigma(f)(a)<c$ on some interval $a\in[-\delta,\delta]_{K}$, and thus $\lVert(\sigma f)(a)\rVert\leq 1$. ∎
We will consider $K$ as a valued field whose valuation ring $K^{\circ}$ is the convex closure of ${\mathbb{R}}$. Denote the maximal ideal of this valuation by $K^{\circ\circ}$, we call elements of $K^{\circ\circ}$ infinitesimal. The corresponding valuation $\lVert\cdot\rVert\colon K^{\times}\mapsto G$ will be called the canonical valuation, and will be written multiplicatively. Note that the residue field of $K$ is isomorphic to ${\mathbb{R}}$, and that the reduction map $K^{\circ}\to K^{\circ}/K^{\circ\circ}$ restricts to an isomorphism on ${\mathbb{R}}\subset K$. For an element $x\in K^{\circ}$, denote by $x^{\circ}$ the element of ${\mathbb{R}}\subset K$ which is closest to $x$, i.e. for which $x-x^{\circ}\in K^{\circ\circ}$. The following lemma tells us that this valuation is automatically henselian. This is again a variation on a classical argument, see e.g. [25, Lem. 2.5], in our more abstract context.
@@
Any ordered field with ${\mathcal{B}}$-analytic structure is henselian with respect to the canonical valuation.
@@
with all $a_{i}\in K^{\circ\circ}$ has a root in $K^{\circ\circ}$ (see e.g. [13, Lem. 2.2]). Replace each $a_{i}$ by a variable $x_{i}$, to obtain $q(x,y)\in{\mathbb{Z}}[x,y]$ such that $q(a,y)=p(y)$, where $a=(a_{2},\dots,a_{d})$. Now consider the polynomial $\tilde{q}(x,y)=q(x-1,y)$ and remark that it satisfies the hypotheses of Lemma 2.1.6. Let $r(x)$ be the implicit function thus defined. Then $(\sigma r)(a)$ is a zero of $p(y-1)$, and moreover $\lVert(\sigma r)(a)\rVert\leq 1$ by Lemma 2.1.8, using that $\lVert r\rVert\leq 1$. ∎
@@
Since the canonical valuation on $K$ is henselian, it extends uniquely to the algebraic closure $K^{\mathrm{alg}}$ of $K$. In particular, this extension coincides on $L$ with the canonical valuation on $L$.
This follows by Weierstrass division, as in e.g. [7, Thm. 2.18]. We sketch the general idea. Take an element $f\in B_{m,\alpha}$ and $a\in L^{m}$, we wish to evaluate $f(a)$. By axiom (1c) and compactness of $[-1,1]_{{\mathbb{R}}}$, it suffices to consider the case where $a\in(L^{\circ\circ})^{m}$. Let $p_{1}(x,y),\dots,p_{m}(x_{m},y)\in{\mathbb{Z}}[x,y]$ with $x=(x_{1},\dots,x_{m})$ and $y=(y_{1},\dots,y_{\ell})$ (for some $\ell\in{\mathbb{N}}$), and $b\in K^{\ell}$ be such that $p_{i}(x_{i},b)$ is the minimal polynomial of $a_{i}$. By henselianity, all conjugates of each $a_{i}$ are infinitesimal in $K^{\mathrm{alg}}$, and hence all $b_{j}$ belong to $K^{\circ\circ}$ for $j=1,\ldots,\ell$. By iterated Weierstrass division of $f(x)$ by $p_{i}(x_{i},y)$ we obtain
@@ 2.2. Almost real closed fields
Almost real closed fields always come equipped with the canonical valuation, for which the valuation ring is the convex closure of ${\mathbb{Q}}$. We will more generally consider convex valuations.
@@
One can show that on an almost real closed field, a valuation is convex if and only if it is henselian, see [20, Lem. 2.1] and [15, Prop. 2.9]. In particular, an ordered field is almost real closed if and only if the canonical valuation is henselian and has real closed residue field.
@@
The canonical valuation has residue field ${\mathbb{R}}$ and is henselian by Lemma 2.1.9. ∎
@@ 2.4. Notation
$K$
an almost real closed field
$K^{\mathrm{rcl}}$
the real closure of $K$
$|\cdot|:K\to K_{>0}$
the absolute value on $K$
$\lVert\cdot\rVert:K\to G$
the canonical valuation on $K$, with multiplicative notation
$K^{\circ},K^{\circ\circ}$
the valuation ring and maximal ideal for the canonical valuation on $K$
$x^{\circ},x\in K^{\circ}$
the element $x^{\circ}$ in ${\mathbb{R}}\subset K$ closest to $x$
$v:K\to\Gamma$
a convex valuation on $K$, with additive notation
${\mathcal{O}}_{K},{\mathcal{M}}_{K}$
the valuation ring and maximal ideal of $v$
${\mathcal{B}}$
a real Weierstrass system over $\Omega$, typically strong and rich
$\Omega$
the value group of the Weierstrass system ${\mathcal{B}}$
@@
the Gauss norm on ${\mathcal{B}}$
${\mathcal{L}}_{{\mathrm{val}}}$
the language of valued fields $\{0,1,+,\cdot,{\mathcal{O}}\}$
${\mathcal{L}}_{\mathrm{or}}$
the language of ordered rings $\{0,1,+,\cdot,<\}$
${\mathcal{L}}_{{\mathcal{B}}}$
the language ${\mathcal{L}}_{\mathrm{or}}$, together with a symbol $(\cdot)^{-1}$ for field inversion, and function symbols for all elements of ${\mathcal{B}}$
${\mathcal{L}}_{\mathrm{ovf}}$
the language of ordered valued fields $\{0,1,+,\cdot,{\mathcal{O}},<\}$
${\mathcal{L}}_{{\mathrm{val}},{\mathcal{B}}}$
the union of the languages ${\mathcal{L}}_{{\mathrm{val}}}$ and ${\mathcal{L}}_{{\mathcal{B}}}$
${\mathcal{L}}_{\mathrm{RV}}$
two-sorted language with ${\mathcal{L}}_{{\mathrm{val}}}$ on the valued field sort and $\{0,1,\cdot,\oplus\}$ on the $\mathrm{RV}$-sort
${\mathcal{L}}_{\mathrm{RV}_{\lambda},{\mathcal{B}}}$
two-sorted language with ${\mathcal{L}}_{{\mathrm{val}},{\mathcal{B}}}$ on the valued field sort and the full induced structure on $\mathrm{RV}_{\lambda}$
${\mathcal{L}}_{{\rm ac},{\mathcal{B}}}$
three-sorted language with symbols for ${\mathcal{B}}$ and an angular component map ${\rm ac}$

### v2

@@ Abstract.
analytic structure, subanalytic sets, quantifier elimination, Weierstrass systems, henselian valued fields, almost real closed fields.
@@ 1. Introduction
Since the work of Łojasiewic, Gabrielov, and Hironaka from the sixties, it has been known that subanalytic sets in ${\mathbb{R}}^{n}$ exhibit tame behaviour similar to the semi-algebraic setting [28, 20, 23, 22]. In more modern terms, the structure ${\mathbb{R}}$ expanded with function symbols for restricted analytic functions is o-minimal. The original works by Łojasiewic, Gabrielov, and Hironaka are rather long and difficult, and work by Denef–van den Dries [14] greatly simplifies the treatment of subanalytic sets through the use of Weierstrass division. Since then, this Weierstrass division has been the central tool for understanding and developing analytic structures on fields, see e.g.  [16, 18, 15, 11, 10, 9, 12, 27, 4]. A comprehensive treatment of fields with analytic structure is the foundational work by Cluckers–Lipshitz [8], which includes both the case of real closed fields and henselian valued fields with analytic structure.
In the current article we continue this line of research, by focussing on fields equipped with real analytic structure, as defined in [8]. Up to now, these structures have only been considered when the field is real closed, leading to o-minimal structures. However, the notion of a field with real analytic structure makes sense even without this assumption, and the aim of this article is to develop the theory of arbitrary fields equipped with real analytic structure. As it turns out, any such field comes with a natural henselian valuation for which the resulting residue field is ${\mathbb{R}}$. In particular, such fields are almost real closed.
@@
Our first main result is that almost real closed fields equipped with ${\mathcal{B}}$-analytic structure exhibit tame behaviour. Of course, in this generality one cannot hope for the theory to be o-minimal or even weakly o-minimal, simply because the field is not necessarily real closed. Instead we show that such fields are $\omega$-h-minimal, a tameness notion for valued fields developed by Cluckers–Halupczok–Rideau [6].
As a consequence of $\omega$-h-minimality, one automatically obtains that definable functions satisfy the Jacobian property, cell decomposition results, and dimension theory. Furthermore, one also obtains bounds on rational points on transcendental definable curves in a suitable sense [5], see Example 5.1.1. This counting of rational points was in fact the original motivation for this project.
@@
Our proof strategy consists of two main steps, and is similar in spirit to that of $\omega$-h-minimality for valued fields with separated analytic structure [6, Thm. 6.2.1]. The first step is a precise analysis of 1-terms, where there is a subtle, yet significant difference compared to [6]. Recall that a main input for [6, Thm. 6.2.1] is a good understanding of 1-terms with parameters in arbitrary models. This follows from an analysis of 1-terms without parameters and a procedure to enlarge the Weierstrass system with constants from the valued field. The latter can be achieved by finding a larger Weierstrass system in which both the valued field and the original Weierstrass system embed, in a compatible way, as in [8]. In our setting, the existence of such a larger system is much less clear, unless our field is a Hahn series equipped with the natural analytic structure (Example 2.1.4).
The Embedding Theorem 3.1.3 is our solution to this problem. Essentially, it allows us to reduce to the nice situation mentioned above, and allows one to develop extension of parameters for real Weierstrass systems. We then build on results from [11] by Cluckers, Lipshitz and Robinson for 1-terms without parameters in real closed fields. Since we are working more generally with fields which are not real closed, this requires some work and care. Since we keep precise track of parameters in each step, we are able to first investigate terms over the real closure and then descend this analysis to the original field. This analysis of terms takes up all of Section 3.
The second step of our proof consists of a relative quantifier elimination statement down to the leading term structure. Here, we follow classical methods by Denef–van den Dries [14] to reduce to the algebraic case, where quantifier elimination relative to the leading term sorts is well-studied (see e.g.  [3, 26, 19]). This is carried out in Section 4.
The notion of $\omega$-h-minimality is really a tameness notion relative to the leading term structure $\mathrm{RV}$. As such, the above theorem does not readily give any information about the induced structure on the residue field and the value group. Therefore, our second main result gives a precise description of the induced structure on the residue field and the value group when $K$ is equipped with its natural valuation. We state the following as an exemplar result, but see Theorem 5.2.1 for more general Weierstrass systems, as defined in Definition 2.1.1. We work in a certain three-sorted language ${\mathcal{L}}_{{\rm ac},{\mathcal{B}}}$ which includes an angular component map ${\rm ac}$, see Section 4.2 for a precise description. Denote by ${\mathbb{R}}_{\mathrm{an}}$ the structure ${\mathbb{R}}$ in the ring language expanded by symbols for restricted analytic functions.
@@
Let ${\mathcal{B}}={\mathcal{A}}((\Omega))$ be the full Weierstrass system over some ordered abelian group $\Omega$, and let $K$ be an almost real closed field with ${\mathcal{B}}$-analytic structure. We consider $K$ as a valued field with its natural valuation. Then
@@
From this relative quantifier elimination, we also obtain an Ax–Kochen–Ersov type result. If $K$ and $K^{\prime}$ are almost real closed fields considered with the natural valuation, then the classical Ax–Kochen–Ersov theorem tells us that $K$ and $K^{\prime}$ are elementarily equivalent as valued fields if and only if their value groups are elementarily equivalent. Indeed, the residue fields are both equal to ${\mathbb{R}}$. The next result can be seen as an extension of this, where we equip $K$ and $K^{\prime}$ with ${\mathcal{B}}$-analytic structure. Denote by ${\mathcal{L}}_{\mathrm{oag}}$ the language of ordered abelian groups $\{0,+,<\}$.
@@
Let ${\mathcal{B}}$ be a real Weierstrass system which is strong and rich, and let $K$ and $K^{\prime}$ be almost real closed fields with ${\mathcal{B}}$-analytic structure. We consider $K$ and $K^{\prime}$ as valued fields with their natural valuation, and denote their value groups by $G$ and $G^{\prime}$. Then
@@ 2.1. Weierstrass systems
We recall the notion of a real Weierstrass system from [8, Def. 3.1.1]. Our definition is in fact slightly different from [8], and the differences will be explained in more detail below.
@@
Consider the system of algebras with $B_{n,\alpha}=A_{n,\alpha}((\Omega))$ for each $\alpha,n$. This is a (rich and strong) real Weierstrass system [8, Rem. 3.13(iv)]. It is called the full Weierstrass system over $\Omega$ and denoted by ${\mathcal{A}}((\Omega))$.
Our definition differs in two ways from [8], so let us indicate the changes. First, axiom 1(e) is new and simply tells us that we can permute the variables of the elements of a Weierstrass system. This seems to be necessary to ensure that composition is well-defined and is also present in the analogous definition of separeted Weierstrass system [9, Def. A.1.1]. Second, in the Weierstrass division axiom the value $\delta\in{\mathbb{R}}_{>0}$ does not depend on $m$, the number of variables of $g$. This will be important for us when expanding Weierstrass systems in Section 3.2. The richness condition is new, and will be needed for us when embedding a field with real analytic structure into a Hahn series field.
The notion of a strong Weierstrass system is also present in [8, Def. 3.1.1] and will be needed in Section 3.4 when analysing terms on the real closure. It roughly allows one to write $f(x,1/x)=g(x)+(1/x)h(1/x)$.
It is important to note that all natural examples of real Weierstrass systems from [8] satisfy the axioms above, in particular they are also real Weierestrass systems according to our definition. Indeed, in the Weierstrass division axiom the value radius $\delta$ depends only on $f$, and not on $g$, as follows from [21, Thm.   II.1, p.  80]. Additionally, all of these examples satisfy the richness condition ((4)). Since our notion of a real Weierstrass system is more restrictive, all results of [8] about them still hold in our context.
@@
This is part of [8, Rem. 3.1.4(ii)] and follows by Weierstrass dividing $1$ by $U(x)$, noting that the latter is regular of degree zero (in any variable). ∎
The following lemma is a real version of the implicit function theorem from [15, Rem. 1.3(4)] and is proved in a similar manner.
@@
We follow the proof of [15, Rem. 1.3(4)]. By Weierstrass division, we may write
@@
The proof idea of the lemma below is classical, see e.g.  [18, Lem. 2.4]. In this more abstract context, some additional input from Lemma 2.1.6 is required, as well as some care about radii of convergence.
@@
Since $\sigma\colon B_{0}\to K$ is order-preserving, (1d) allows us to reduce to the case that $\lVert f\rVert=1$. Using axiom (1c) and compactness of $[-1,1]_{{\mathbb{R}}}$, it suffices to prove this lemma for $a\in[-\delta,\delta]_{K}$ for some $\delta\leq\alpha$ with $\delta\in{\mathbb{R}}_{>0}$. Now, there exist some $b,c\in{\mathbb{R}}$ such that $b<f_{0}(0)<c$. By Lemma 2.1.7, applied to $f(x)-b$ and $f(x)-c$, it follows that also $b<\sigma(f)(a)<c$ on some interval $a\in[-\delta,\delta]_{K}$, and thus $\lVert(\sigma f)(a)\rVert=1$. ∎
We consider $K$ as a valued field whose valuation ring $K^{\circ}$ is the convex closure of ${\mathbb{R}}$. Denote the maximal ideal of this valuation by $K^{\circ\circ}$, we call elements of $K^{\circ\circ}$ infinitesimal. The corresponding valuation $\lVert\cdot\rVert\colon K^{\times}\mapsto G$ will be called the natural valuation, and will be written multiplicatively. Note that the residue field of $K$ is isomorphic to ${\mathbb{R}}$, and that the reduction map $K^{\circ}\to K^{\circ}/K^{\circ\circ}$ restricts to an isomorphism on ${\mathbb{R}}\subset K$. For an element $x\in K^{\circ}$, denote by $x^{\circ}$ the element of ${\mathbb{R}}\subset K$ which is closest to $x$, i.e. for which $x-x^{\circ}\in K^{\circ\circ}$. The following lemma tells us that this valuation is automatically henselian. This is again a variation on a classical argument, see e.g. [18, Lem. 2.5], in our more abstract context.
@@
Any ordered field with ${\mathcal{B}}$-analytic structure is henselian with respect to the natural valuation.
@@
with all $a_{i}\in K^{\circ\circ}$ has a root in $K^{\circ\circ}$ (see e.g. [17, Lem. 2.2]). Replace each $a_{i}$ by a variable $x_{i}$, to obtain $q(x,y)\in{\mathbb{Z}}[x,y]$ such that $q(a,y)=p(y)$, where $a=(a_{2},\dots,a_{d})$. Now consider the polynomial $\tilde{q}(x,y)=q(x-1,y)$ and remark that it satisfies the hypotheses of Lemma 2.1.6. Let $r(x)$ be the implicit function thus defined. Then $(\sigma r)(a)$ is a zero of $p(y-1)$, and moreover $\lVert(\sigma r)(a)\rVert\leq 1$ by Lemma 2.1.8, using that $\lVert r\rVert\leq 1$. ∎
@@
Since the natural valuation on $K$ is henselian, it extends uniquely to the algebraic closure $K^{\mathrm{alg}}$ of $K$. In particular, this extension coincides on $L$ with the natural valuation on $L$.
This follows by Weierstrass division, as in e.g. [10, Thm. 2.18]. We sketch the general idea. Take an element $f\in B_{m,\alpha}$ and $a\in L^{m}$, we wish to evaluate $f(a)$. By axiom (1c) and compactness of $[-1,1]_{{\mathbb{R}}}$, it suffices to consider the case where $a\in(L^{\circ\circ})^{m}$. Let $p_{1}(x,y),\dots,p_{m}(x_{m},y)\in{\mathbb{Z}}[x,y]$ with $x=(x_{1},\dots,x_{m})$ and $y=(y_{1},\dots,y_{\ell})$ (for some $\ell\in{\mathbb{N}}$), and $b\in K^{\ell}$ be such that $p_{i}(x_{i},b)$ is the minimal polynomial of $a_{i}$. By henselianity, all conjugates of each $a_{i}$ are infinitesimal in $K^{\mathrm{alg}}$, and hence all $b_{j}$ belong to $K^{\circ\circ}$ for $j=1,\ldots,\ell$. By iterated Weierstrass division of $f(x)$ by $p_{i}(x_{i},y)$ we obtain
@@ 2.2. Almost real closed fields
Almost real closed fields always come equipped with the natural valuation, for which the valuation ring is the convex closure of ${\mathbb{Q}}$. We will more generally consider convex valuations.
@@
One can show that on an almost real closed field, a valuation is convex if and only if it is henselian, see [24, Lem. 2.1] and [13, Prop. 2.9]. In particular, an ordered field is almost real closed if and only if the natural valuation is henselian and has real closed residue field.
@@
The natural valuation has residue field ${\mathbb{R}}$ and is henselian by Lemma 2.1.9. ∎
@@ 2.4. Notation
$K$ an almost real closed field
$K^{\mathrm{rcl}}$ the real closure of $K$
$|\cdot|:K\to K_{>0}$ the absolute value on $K$
$\lVert\cdot\rVert:K\to G$ the natural valuation on $K$, with multiplicative notation
$K^{\circ},K^{\circ\circ}$ the valuation ring and maximal ideal for the natural valuation on $K$
$x^{\circ},x\in K^{\circ}$ the element $x^{\circ}$ in ${\mathbb{R}}\subset K$ closest to $x$
$v:K\to\Gamma$ a convex valuation on $K$, with additive notation
${\mathcal{O}}_{K},{\mathcal{M}}_{K}$ the valuation ring and maximal ideal of $v$
${\mathcal{B}}$ a real Weierstrass system over $\Omega$, typically strong and rich
$\Omega$ the value group of the Weierstrass system ${\mathcal{B}}$
@@
${\mathcal{L}}_{{\mathrm{val}}}$ the language of valued fields $\{0,1,+,\cdot,{\mathcal{O}}\}$
${\mathcal{L}}_{\mathrm{or}}$ the language of ordered rings $\{0,1,+,\cdot,<\}$
${\mathcal{L}}_{{\mathcal{B}}}$ the language ${\mathcal{L}}_{\mathrm{or}}$, together with a symbol $(\cdot)^{-1}$ for field inversion, and function symbols for all elements of ${\mathcal{B}}$
${\mathcal{L}}_{\mathrm{ovf}}$ the language of ordered valued fields $\{0,1,+,\cdot,{\mathcal{O}},<\}$
${\mathcal{L}}_{{\mathrm{val}},{\mathcal{B}}}$ the union of the languages ${\mathcal{L}}_{{\mathrm{val}}}$ and ${\mathcal{L}}_{{\mathcal{B}}}$
${\mathcal{L}}_{\mathrm{RV}}$ two-sorted language with ${\mathcal{L}}_{{\mathrm{val}}}$ on the valued field sort and $\{0,1,\cdot,\oplus\}$ on the $\mathrm{RV}$-sort
${\mathcal{L}}_{\mathrm{RV}_{\lambda},{\mathcal{B}}}$ two-sorted language with ${\mathcal{L}}_{{\mathrm{val}},{\mathcal{B}}}$ on the valued field sort and the full induced structure on $\mathrm{RV}_{\lambda}$
${\mathcal{L}}_{{\rm ac},{\mathcal{B}}}$ three-sorted language with symbols for ${\mathcal{B}}$ and an angular component map ${\rm ac}$